Please help. y = 2a - ab + 3, for a.
Appreciate any help.
Thanks.

Answer by jb007mi6(53)

so you have to isolate the "a"
switch the equation around for mathematical purposes
2a-ab+3=y
subtract both sides by 3
2a-ab+3-3=y-3
this is
2a-ab=y-3
factor out an "a"
a(2-b)=y-3
divide both sides by "2-b" to isolate the "a"
a=(y-3)/(2-b)